# Autodeleveraging: Impossibilities and Optimization

### v1

@@ F.4 Follower Strategic Responses
F.4.1 Adverse Selection Under Pro-Rata
@@
Lemma F.7 (When RAP burdens a trader less than Pro-Rata).
@@
Corollary F.8 (Pro-Rata death spiral vs. RAP retention).
Fix a type $i$ and suppose there is a set of rounds of positive probability on which $\lambda_{t,i}g(\lambda_{t,i})\leq\mu^{(g)}_{t}$ (so $s^{\mathrm{RAP}}_{t,i}\leq s^{\mathrm{PR}}_{t,i}$ by Lemma F.7). If, over the same distribution of rounds,
@@
RAP with $g(\lambda)=\lambda$. Weights scale as $e_{i}\lambda_{i}^{2}$, so $L$’s share collapses to $s_{L}^{\mathrm{RAP}}=240/(240+1440)\approx 0.14$ and $H$’s share rises to $0.86$. The round 1 haircut factors become $h_{L}^{\mathrm{RAP}}\approx 0.095$ and $h_{H}^{\mathrm{RAP}}\approx 0.86$ (masses $H^{\mathrm{RAP}}=(5.7,34.3)$), yielding $U^{\mathrm{RAP}}=(6.3,5.7)>u_{0}$, so both types remain for round 2. With both accounts active in round 2, shares remain tilted toward $H$ and haircuts $(4.3,25.7)$ keep both traders above $u_{0}$, preventing churn and stabilizing $W_{t}$. This concrete two-player, two-round example mirrors the equilibria predicted by Appendix B.6 and the empirical replay in §9: pro-rata drives the safest capital away first, whereas RAP reallocates the burden toward high-leverage accounts and keeps the equity base intact.
F.4.2 Waiting Game and MDIC-NW
@@
When an ADL event creates a deficit $D_{t}$, the exchange moves first: it announces the contemporaneous severity $\theta_{t}$ (hence the haircut budget $\theta_{t}D_{t}$) together with the liquidity premium $\kappa_{t}$ it is willing to pay per unit of external capital. [^14] In practice, this occurs via either an increase in the expected payment for users who stake to an insurance fund (e.g. [DriftInsuranceFund]) or add assets to an HLP/LLP style vault [HyperliquidHLPVaults, LighterWhitepaper] A Backstop Liquidity Provider (BLP) then decides whether to intervene immediately by injecting any $q_{t}\in[0,D_{t}]$ units, or to wait $u\geq 1$ additional rounds before posting the same quantity. Waiting exposes the BLP to time discounting (with a factor $\beta^{u}$) and to the future premium schedule $(\theta_{t+u},\kappa_{t+u})$. The BLP’s payoff from intervening in round $t+u$ with size $q$ is $\beta^{u}q(\theta_{t+u}-\kappa_{t+u})$, while the exchange’s objective is to restore solvency before a deadline by ensuring that the entire deficit is filled (so delay harms solvency).
This Stackelberg game involves the exchange posting incentives as leader and BLPs optimally choose a stopping time and quantity as a follower. The setup captures the “waiting game” intuition: unless the contract guarantees non-negative per-unit surplus right now, rational liquidity providers will defer, pushing resolution beyond the solvency window. We note that this resembles other waiting games in Maximal Extractable Value (MEV) auctions that have been studied empirically on live systems [messias2025express, mamageishvili2025timeboost].
@@
Lemma F.9 (Per-unit No-Wait implies immediate action).
@@
Corollary F.10 (No-Wait bounds solvency recovery).
@@
Lemma F.9 implies that at $\tau_{\mathrm{def}}$ the BLP injects $q_{\tau_{\mathrm{def}}}^{\star}=D_{\tau_{\mathrm{def}}}$, so the entire deficit is covered immediately. Therefore the insurance fund (or deficit buffer) reaches the safety level $\delta$ after the same round, yielding $\tau_{\mathrm{solv}}=\tau_{\mathrm{def}}+1$. ∎
@@
Let $U(\tau,q)=\beta^{\tau-t}q(\theta_{\tau}-\kappa_{\tau})$ be the discounted utility of a BLP who provides liquidity $q$ at time $\tau\geq t$. Here, $\theta_{\tau}$ represents the payment per unit of liquidity (derived from the ADL severity) and $\kappa_{\tau}$ represents the cost per unit (liquidity premium). The BLP solves the optimal stopping problem (see, e.g. [Peskir2006]):
@@
Appendix G Price of Anarchy Phase Transitions
@@
G.1 Fairness Phase Transition
@@ G.1.1 Welfare and Assumptions
$W_{\mathrm{Fair}}(\pi)\;:=\;\mathop{\bf E{}}\left[\frac{\omega_{T}^{\pi}}{D_{T}^{\pi}}\right].$
This metric captures the efficiency of haircuts: higher values imply that the policy covers deficits $D_{T}^{\pi}$ while preserving maximal winner equity $\omega_{T}^{\pi}$. Extreme-value scaling implies $W_{\mathrm{Fair}}$ scales inversely with the severity load.
@@
We require the following regularity assumptions. LLN and EV Scaling (Prop. G.1) establish the baseline scales for deficits ($O(n)$) and equity ($O(b_{n})$). Anti-concentration (Prop. G.2) ensures that $b_{n}\ll n$, creating the scarcity of winner equity that drives the phase transition.
Proposition G.1 (LLN and EV scaling).
@@
Proposition G.2 (Anti-concentration).
@@ G.1.2 Phase Transition
Theorem G.3 (Fairness PoA Phase Transition).
Suppose Assumptions G.1G.2 hold. Let $\pi^{\star}$ be the optimal dynamic policy and $\pi^{\mathrm{stat}}$ be any static policy with load $\kappa_{n}^{\mathrm{stat}}$.
@@
First, for the dynamic benchmark, the optimal Stackelberg policy $\pi^{\star}$ minimizes haircuts ex-post to cover the realized deficit $D_{T}\asymp n$. By targeting the haircut to preserve the top winner’s equity $\omega_{T}\asymp b_{n}$, the dynamic policy achieves a Profitability-to-Total-Solvency Ratio (PTSR) that concentrates around a constant, $W_{\mathrm{Fair}}(\pi^{\star})\asymp 1$ (Theorem B.3). This is possible because the controller can observe the realization of heavy-tailed variates and adjust the severity $\theta$ precisely to the minimal necessary level.
Second, for the static policy, the severity $\theta_{n}$ is fixed ex-ante and applies a uniform pressure $\theta_{n}n$ on the book, which must be absorbed by individual winners with capacity scaling as $b_{n}$. Under heavy-tailed scaling, Proposition 5.1 shows that this mismatch leads to a welfare decay $W_{\mathrm{Fair}}(\pi)\asymp b_{n}/(\theta_{n}n)=1/\kappa_{n}$, as the policy blindly destroys small winners or fails to extract sufficient capital from the tail without excessive rates. Crucially, as $n\to\infty$, the gap between the aggregate load $O(n)$ and individual capacity $O(b_{n})$ widens (since $b_{n}=o(n)$), forcing $\kappa_{n}$ to grow if $\theta_{n}$ does not vanish rapidly enough.
@@
Example G.4 (Light-tailed Failure).
@@
G.2 Revenue Phase Transition
@@ G.2.1 Joint Welfare
G.2.2 Phase Transition
Proposition G.5 (Revenue PoA Phase Transition).
@@ G.2.2 Phase Transition
Second, a static policy is parameterized by a fixed fee diversion rate $\delta\in[0,1]$ chosen ex-ante, meaning it diverts a constant fraction of fees $\mathcal{D}_{t}=\delta\phi_{t}$ (where $\Phi_{T}=\sum\phi_{t}$ follows the scaling in Assumption I.2) to the insurance fund. To build a fund $K_{t}$ capable of absorbing heavy-tailed shocks with infinite variance ($\alpha<2$), Standard ruin theory [AsmussenAlbrecher2010, Embrechts1997] implies that to ensure Solvency Condition (S) (Definition I.4) against the maximum jump $\Delta_{T}\sim n^{1/\alpha}$, the policy must essentially set $\delta\to 1$ to handle the timing mismatch where large shocks occur before the fund accumulates. Any policy that attempts to maintain $\delta<1$ will fail solvency with probability approaching 1, while a policy with $\delta\approx 1$ consumes the entire revenue stream, driving $W_{\mathrm{Rev}}(\pi^{\mathrm{stat}})\to 0$ or into negative territory due to insolvency penalties.
@@ G.3 The Aggregation Paradox
Proposition G.6 (Sum vs. Min Aggregation).
Let $W_{\Sigma}=W_{\mathrm{Fair}}+W_{\mathrm{Rev}}$ and $W_{\min}=\min(W_{\mathrm{Fair}},W_{\mathrm{Rev}})$ be normalized combined objectives.
@@
    Sum-Welfare is Bounded: $\mathrm{PoA}_{\Sigma}\leq 2$. A static policy can always choose to satisfy one objective fully, achieving at least half the optimal total score.
@@
    Min-Welfare is Unbounded: $\mathrm{PoA}_{\min}\to\infty$. In the heavy-tailed regime, static policies face the ADL Trilemma and must drive at least one objective to zero, while dynamic policies maintain both.

### v3

@@ G.4 Follower Strategic Responses
G.4.1 Adverse Selection Under Pro-Rata
@@
Lemma G.7 (When RAP burdens a trader less than Pro-Rata).
@@
Corollary G.8 (Pro-Rata death spiral vs. RAP retention).
Fix a type $i$ and suppose there is a set of rounds of positive probability on which $\lambda_{t,i}g(\lambda_{t,i})\leq\mu^{(g)}_{t}$ (so $s^{\mathrm{RAP}}_{t,i}\leq s^{\mathrm{PR}}_{t,i}$ by Lemma G.7). If, over the same distribution of rounds,
@@
RAP with $g(\lambda)=\lambda$. Weights scale as $e_{i}\lambda_{i}^{2}$, so $L$’s share collapses to $s_{L}^{\mathrm{RAP}}=240/(240+1440)\approx 0.14$ and $H$’s share rises to $0.86$. The round 1 haircut factors become $h_{L}^{\mathrm{RAP}}\approx 0.095$ and $h_{H}^{\mathrm{RAP}}\approx 0.86$ (masses $H^{\mathrm{RAP}}=(5.7,34.3)$), yielding $U^{\mathrm{RAP}}=(6.3,5.7)>u_{0}$, so both types remain for round 2. With both accounts active in round 2, shares remain tilted toward $H$ and haircuts $(4.3,25.7)$ keep both traders above $u_{0}$, preventing churn and stabilizing $W_{t}$. This concrete two-player, two-round example mirrors the equilibria predicted by Appendix C.6 and the empirical replay in §9: pro-rata drives the safest capital away first, whereas RAP reallocates the burden toward high-leverage accounts and keeps the equity base intact.
G.4.2 Waiting Game and MDIC-NW
@@
When an ADL event creates a deficit $D_{t}$, the exchange moves first: it announces the contemporaneous severity $\theta_{t}$ (hence the haircut budget $\theta_{t}D_{t}$) together with the liquidity premium $\kappa_{t}$ it is willing to pay per unit of external capital. [^16] In practice, this occurs via either an increase in the expected payment for users who stake to an insurance fund (e.g. [48]) or add assets to an HLP/LLP style vault [71, 82] A Backstop Liquidity Provider (BLP) then decides whether to intervene immediately by injecting any $q_{t}\in[0,D_{t}]$ units, or to wait $u\geq 1$ additional rounds before posting the same quantity. Waiting exposes the BLP to time discounting (with a factor $\beta^{u}$) and to the future premium schedule $(\theta_{t+u},\kappa_{t+u})$. The BLP’s payoff from intervening in round $t+u$ with size $q$ is $\beta^{u}q(\theta_{t+u}-\kappa_{t+u})$, while the exchange’s objective is to restore solvency before a deadline by ensuring that the entire deficit is filled (so delay harms solvency).
This Stackelberg game involves the exchange posting incentives as leader and BLPs optimally choose a stopping time and quantity as a follower. The setup captures the “waiting game” intuition: unless the contract guarantees non-negative per-unit surplus right now, rational liquidity providers will defer, pushing resolution beyond the solvency window. We note that this resembles other waiting games in Maximal Extractable Value (MEV) auctions that have been studied empirically on live systems [86, 83].
@@
Lemma G.9 (Per-unit No-Wait implies immediate action).
@@
Corollary G.10 (No-Wait bounds solvency recovery).
@@
Lemma G.9 implies that at $\tau_{\mathrm{def}}$ the BLP injects $q_{\tau_{\mathrm{def}}}^{\star}=D_{\tau_{\mathrm{def}}}$, so the entire deficit is covered immediately. Therefore the insurance fund (or deficit buffer) reaches the safety level $\delta$ after the same round, yielding $\tau_{\mathrm{solv}}=\tau_{\mathrm{def}}+1$. ∎
@@
Let $U(\tau,q)=\beta^{\tau-t}q(\theta_{\tau}-\kappa_{\tau})$ be the discounted utility of a BLP who provides liquidity $q$ at time $\tau\geq t$. Here, $\theta_{\tau}$ represents the payment per unit of liquidity (derived from the ADL severity) and $\kappa_{\tau}$ represents the cost per unit (liquidity premium). The BLP solves the optimal stopping problem (see, e.g. [95]):
@@
Appendix H Price of Anarchy Phase Transitions
@@
H.1 Fairness Phase Transition
@@ H.1.1 Welfare and Assumptions
$W_{\mathrm{Fair}}(\pi)\;:=\;\mathop{\bf E{}}\left[\frac{\upsilon_{T}^{\pi}}{D_{T}^{\pi}}\right],$
where $\upsilon_{T}^{\pi}$ is the maximum post-ADL endowment (under PNL-only, maximum post-ADL positive PNL). This metric captures the efficiency of haircuts: higher values imply that the policy covers deficits $D_{T}^{\pi}$ while preserving maximal winner endowment $\upsilon_{T}^{\pi}$. Extreme-value scaling implies $W_{\mathrm{Fair}}$ scales inversely with the severity load.
@@
We require the following regularity assumptions. LLN and EV Scaling (Prop. H.1) establish the baseline scales for deficits ($O(n)$) and equity ($O(b_{n})$). Anti-concentration (Prop. H.2) ensures that $b_{n}\ll n$, creating the scarcity of winner equity that drives the phase transition.
Proposition H.1 (LLN and EV scaling).
@@
Proposition H.2 (Anti-concentration).
@@ H.1.2 Phase Transition
Theorem H.3 (Fairness PoA Phase Transition).
Suppose Assumptions H.1H.2 hold. Let $\pi^{\star}$ be the optimal dynamic policy and $\pi^{\mathrm{stat}}$ be any static policy with load $\kappa_{n}^{\mathrm{stat}}$.
@@
First, for the dynamic benchmark, the optimal Stackelberg policy $\pi^{\star}$ minimizes haircuts ex-post to cover the realized deficit $D_{T}\asymp n$. By targeting the haircut to preserve the top winner’s endowment $\upsilon_{T}\asymp b_{n}$ (under PNL-only, maximum positive PNL), the dynamic policy achieves a Profitability-to-Total-Solvency Ratio (PTSR) that concentrates around a constant, $W_{\mathrm{Fair}}(\pi^{\star})\asymp 1$ (Theorem C.3). This is possible because the controller can observe the realization of heavy-tailed variates and adjust the severity $\theta$ precisely to the minimal necessary level.
Second, for the static policy, the severity $\theta_{n}$ is fixed ex-ante and applies a uniform pressure $\theta_{n}n$ on the book, which must be absorbed by individual winners with endowment capacity scaling as $b_{n}$. Under heavy-tailed scaling, Proposition 5.1 shows that this mismatch leads to a welfare decay $W_{\mathrm{Fair}}(\pi)\asymp b_{n}/(\theta_{n}n)=1/\kappa_{n}$, as the policy blindly destroys small winners’ endowments or fails to extract sufficient capital from the tail without excessive rates. Crucially, as $n\to\infty$, the gap between the aggregate load $O(n)$ and individual capacity $O(b_{n})$ widens (since $b_{n}=o(n)$), forcing $\kappa_{n}$ to grow if $\theta_{n}$ does not vanish rapidly enough.
@@
Example H.4 (Light-tailed Failure).
@@
H.2 Revenue Phase Transition
@@ H.2.1 Joint Welfare
H.2.2 Phase Transition
Proposition H.5 (Revenue PoA Phase Transition).
@@ H.2.2 Phase Transition
Second, a static policy is parameterized by a fixed fee diversion rate $\delta\in[0,1]$ chosen ex-ante, meaning it diverts a constant fraction of fees $\mathcal{D}_{t}=\delta\phi_{t}$ (where $\Phi_{T}=\sum\phi_{t}$ follows the scaling in Assumption J.2) to the insurance fund. To build a fund $K_{t}$ capable of absorbing heavy-tailed shocks with infinite variance ($\alpha<2$), Standard ruin theory [16, 52] implies that to ensure Solvency Condition (S) (Definition J.4) against the maximum jump $\Delta_{T}\sim n^{1/\alpha}$, the policy must essentially set $\delta\to 1$ to handle the timing mismatch where large shocks occur before the fund accumulates. Any policy that attempts to maintain $\delta<1$ will fail solvency with probability approaching 1, while a policy with $\delta\approx 1$ consumes the entire revenue stream, driving $W_{\mathrm{Rev}}(\pi^{\mathrm{stat}})\to 0$ or into negative territory due to insolvency penalties.
@@ H.3 The Aggregation Paradox
Proposition H.6 (Sum vs. Min Aggregation).
Let the normalized combined objectives be
$W_{\Sigma}=W_{\mathrm{Fair}}+W_{\mathrm{Rev}},\qquad W_{\min}=\min(W_{\mathrm{Fair}},W_{\mathrm{Rev}}).$
@@
    Sum-Welfare is Bounded. A static policy can always choose to satisfy one objective fully, achieving at least half the optimal total score; consequently
    $\mathrm{PoA}_{\Sigma}\leq 2.$
@@
    Min-Welfare is Unbounded. In the heavy-tailed regime, static policies face the ADL Trilemma and must drive at least one objective to zero, while dynamic policies maintain both, so
    $\mathrm{PoA}_{\min}\to\infty.$